A pub uses a periodic review system where it checks the liquor stock every Tuesday and place an order which arrives on Thursday. The pub has observed that demand for beer averages 7 units per day with a standard deviation of 2 units. In any given week, they would like to be 90% sure that they don't run out of beer before the order arrives. On Tuesday, if they have 10 units of beer on hand, how many units should they order?


a. 56.84 units


b. 59.00 units c. 41.70 units


d. 52.60 units


1
Expert's answer
2021-04-25T09:07:20-0400

"\mu=7,\sigma=2"

"\alpha=0.1"


The value of bottle used"X=\mu+Z_{0.05}\dfrac{\sigma}{\sqrt{n}}" ="7+2.235\times \dfrac{2}{\sqrt{100}}=7+0.447=7.447"


Total number of beers used from Tuesday to next tuesday "= 7.447\times6+10=54.68"


Hence The number of units they have to order is 54.68.